two cars are 12 miles apart and are approaching each other on a straight road. each car is moving at 60 mi/h relative to the road. how long will it be before the cars pass each other? give your answer in minutes
1
Expert's answer
2019-10-31T11:18:10-0400

Let "d = 12 mi", "v = 60\frac{mi}{h}".

Let the initial positions of the cars be "x_0" and "x_0 + d" respectively, and the first car move in the positive "ox" direction, while the other car in the negative. Hence, the equations of motion of the cars are "x_1(t) = x_0 + v t", "x_2(t) = (x_0 + d) - v t". When they pass each other, "x_1(t) = x_2(t)", from where "2 vt = d" and "t = \frac{d}{2 v} \approx 0.1 h = 6 min".